OpenStudy (shayaan_mustafa):

Understanding how to find intervals, where functions 1)increase and decrease using 1st Derivative Test. 2)concave up and concave down using 2nd Derivative Test. Question is; \[f(x)=\sqrt[3]{x+2}\]

OpenStudy (shayaan_mustafa):

I know where F'(x)>0 then increase and f'(x)<0decrease. similarly f''(x)>0 then concave up and f''(x)<0 then concave down. But is there any specific procedure to determine these properties?

OpenStudy (anonymous):

you don't really need calc for the first one. \[f(x)=\sqrt[3]{x+2}\] looks just like \[\sqrt[3]{x}\] shifted to the left two units. it is always increasing

OpenStudy (anonymous):

the bigger the number, the larger the cubed root

OpenStudy (shayaan_mustafa):

ok. if x is negative then?

OpenStudy (anonymous):

\[a<b\implies \sqrt[3]{a}<\sqrt[3]{b}\] i

OpenStudy (shayaan_mustafa):

I know what are you saying. But there is also a specific technique to find increasing and decreasing function by differentiation.

OpenStudy (amistre64):

step a, find the places where the derivaite (the slope of any point on the curve) is zero

OpenStudy (shayaan_mustafa):

at x=-2

OpenStudy (amistre64):

step 2, pick a point inbetween zeros to test for sign

OpenStudy (amistre64):

when sign of the derivative is negative; you are decreasing; when its positive you are increasing; when its zero you are neither increasing nor decreasing

OpenStudy (amistre64):



OpenStudy (shayaan_mustafa):

But this function will never be decreasing.

OpenStudy (amistre64):

which is why its always increasing except at f'(x) = 0

OpenStudy (amistre64):

the same idea with the second derivative to determine cave up or cave down

OpenStudy (shayaan_mustafa):

\[\LARGE f(x)=\frac{1}{3(x+2)^{\frac{2}{3}}}\] It will never show - sign.

OpenStudy (amistre64):

find all critical values; zeros AND undefineds

OpenStudy (shayaan_mustafa):

Sorry it is f'(x)

OpenStudy (amistre64):

at x=-2 we are undefined; so test both sides of x=-2

OpenStudy (amistre64):

and your right, its never goes negative so that tells us?

OpenStudy (shayaan_mustafa):

One thing I want to ask here. Do for both f '(x) and f ''(x) we have to find critical points?

OpenStudy (shayaan_mustafa):

It tells us that it will always be increasing function.

OpenStudy (amistre64):

yes, the critical points partition the interval into testable portions

OpenStudy (shayaan_mustafa):

(-inf , +inf) increasing interval.

OpenStudy (amistre64):

good, and since it never goes zero we can determine that x=-2 is also an increasing spot

OpenStudy (amistre64):

or do we leave that undefined?

OpenStudy (amistre64):

the slope is vertical at x=-2 ...

OpenStudy (amistre64):

i forget what do do with x=-2 :) it might need to be opened up

OpenStudy (shayaan_mustafa):

\[\LARGE f''(x)=\frac{-2}{4(x+2)^\frac{5}{3}}\] It will be negative if x>-2 and positive if x<-2 So concave up interval is (-inf, -2) And concave down interval is (-2, +inf) Am I right?

OpenStudy (shayaan_mustafa):

Sorry it is 9 in the denominator not 4

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=derivative+cbrt%28x%2B2%29 this is showing something wierd to me

OpenStudy (shayaan_mustafa):

Yes it is first derivative. and I have showed that is 2nd derivative.

OpenStudy (amistre64):

why is the wolf showing a negative portion on the left of x=-2?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3Dcbrt%28x%2B2%29 and this is wacky too; its not showing a translation if cbrt(x)

OpenStudy (shayaan_mustafa):

Firstly I want to tell you one thing which is more important. I have never used any type of graphing utility or any calculator in my life. So I don't know how wolfram alpha works and what it shows. But I believe upon my calculations by hand and mind. :)

OpenStudy (amistre64):

oh i trust youre correct since I know what cbrt(x+2) is spose to look like. but the wolf is usually a good dbl chk to it

OpenStudy (amistre64):

I just cant remember what to do at x=-2 with an undefined slope; if we include it in the interval or not

OpenStudy (amistre64):

id say exclude it, since a vertical slope is neither inc or dec ....

OpenStudy (shayaan_mustafa):

Now One thing is clear to me. I have learned how to check increament and decrement of the function. But want to know about concavity.

OpenStudy (amistre64):

concavity is the same thing; but with f''

OpenStudy (shayaan_mustafa):

Howard Anton Calculus says to include that part i.e. -2 in our case. But some where I have seen not to include critical point. I just got confused :(

OpenStudy (amistre64):

id say exclude it; vertical and horizontals dont seem to be defined as inc or dec

OpenStudy (amistre64):

but at any rate, use the material that you are using to define it :)

OpenStudy (shayaan_mustafa):

As for f ' (x) we have find critical point. Do for f '' (x) we have to find critical point separately in order to test for sign? Can you get what I am trying to say?

OpenStudy (amistre64):

yes, regards f'' as a separate beast to play with

OpenStudy (amistre64):

if it happens to have the same critical points, so be it; but never assume it :)

OpenStudy (shayaan_mustafa):

Yeah i am using howard anton calculus. famous calculus in the world. so I should include critical point. ok i got you.. :) Yes. I got what you have suggested to me.

OpenStudy (shayaan_mustafa):

Now one last thing. what about inflection point?

OpenStudy (amistre64):

inflection is where we change concavity; the critical points that have opposite sides to them are inflection; usually where f''(x) = 0 or is undefined need to be considered

OpenStudy (shayaan_mustafa):

I know where curve changes its concavity that is inflection point. but which critical point we have to take as a inflection point? That one which we have find for f ' (x) or that one which we have find for f '' (x)?

OpenStudy (shayaan_mustafa):

What do you mean by this "that have opposite sides to them are inflection"

OpenStudy (amistre64):



OpenStudy (amistre64):

where c is a critical point

OpenStudy (shayaan_mustafa):

OK. I have got all that. Thanks once again. :) Be happy always. :)

OpenStudy (amistre64):

good luck :)